OpenStudy (tinalikescats):

Which of the following is the correct expansion of the binomial (x + y)6? x6 + 6x5y + 15x4y2 + 20x3y3 + 15x2y4 + 6xy5 + y6 x6 + 155y + 6x4y2 + 20x3y3 + 6x2y4 + 15xy5 + y6 x6 + 6x5y5 + 15x4y4 + 20x3y3 + 15x2y2 + 6xy1 + y6 x6 + 6x5y + 20x4y2 + 15x3y3 + 20x2y4 + 6xy5 + y6

OpenStudy (john_es):

Do you remember the binomial expansion?

OpenStudy (john_es):

And do you know the Pascal triangle?

OpenStudy (tinalikescats):

uh no i dont

OpenStudy (tinalikescats):

like i know it but i don't remember it

OpenStudy (john_es):

Ok. The binomial expansion tell you this, for example, for the the 2 and 3 powers, \[(x+y)^2=x^2+2xy+y^2\] \[(x+y)^3=x^3+3x^2y+3xy^2+y^3\]

OpenStudy (tinalikescats):

what

OpenStudy (john_es):

Do you see how it goes?

OpenStudy (tinalikescats):

no :(

OpenStudy (john_es):

You can do the same thing with the sixth power, \[(x+y)^6=x^6+6x^5y+15x^4y^2+20x^3y^3+15x^2y^4+6xy^5+y^6\]

OpenStudy (anonymous):



OpenStudy (tinalikescats):



OpenStudy (john_es):

The way to obtain this is just to multply, as @BPDlkeme567 says, \[(x+y)(x+y)(x+y)(x+y)(x+y)(x+y)\]

OpenStudy (anonymous):

can you not multiply (x + y)(x+y)?

OpenStudy (tinalikescats):

no ;-;

OpenStudy (anonymous):

Okay so we need to start there

OpenStudy (tinalikescats):

c:

OpenStudy (john_es):

Yes, it is better the way @BPDlkeme567 says.

OpenStudy (anonymous):

that is known as the difference of two squares. i.e x^2 + y^2

OpenStudy (anonymous):

so, (x + y)(x- -y) = x^2 + y^2

OpenStudy (anonymous):

hey human calculator @John_ES , can you take over and teach binomial theorem?

OpenStudy (tinalikescats):

and polynomial identities

OpenStudy (anonymous):

@tinalikescats it is actually based on a very simple idea, invented by the chines (forget that people call it Pascal's triangle), but have a look on wikipedia at Pascal's triangle.

OpenStudy (tinalikescats):

wot

OpenStudy (anonymous):

Pascals triangle starts with 1, next two numbers must be put below this, which is the sum of the previous numbers, but since there is only 1 previous number the two numbers under this are 1

OpenStudy (anonymous):

oh, sorry, this might seem irrelevant, but these are the co-efficients of the mathematical expressions

OpenStudy (anonymous):

are you there?

OpenStudy (mathexpert77):

x6 + 6x5y + 15x4y2 + 20x3y3 + 15x2y4 + 6xy5 + y6